# Similarity-Navigated Conformal Prediction for Graph Neural Networks

### v1

@@ 4.2 Experimental results
SNAPS generates smaller prediction sets and achieves higher singleton hit ratio. Table 1 shows that $\mathrm{Coverage}$ of all conformal prediction methods is close to the desired coverage $1-\alpha$. At a significance level $\alpha=0.05$, $\mathrm{Size}$ and $\mathrm{SH}$ exhibit superior performance. For example, when evaluated on Products, SNAPS reduces $\mathrm{Size}$ from 14.92 of APS to 7.68. Overall, the experiments show that SNAPS has the desired coverage rate and gets smaller $\mathrm{Size}$ and higher $\mathrm{SH}$ than APS, RAPS and DAPS. Detailed results for other basic models and SNAPS based on RAPS are available in Appendix C.
SNAPS generates smaller average prediction sets for each label. We conduct additional experiments to analyse the average performance of APS and SNAPS on nodes belonging to the same label at a significance level $\alpha=0.05$. Figure LABEL:fig:mean-size-aps shows that the distribution of the average non-conformity scores for nodes belonging to the same label aligns with the assumptions made in Proposition 2, i.e., $E_{k}[\boldsymbol{S}_{uk}]\ll\eta$ and $E_{k}[\boldsymbol{S}_{ui}]-\eta\geq-\Delta$, where $\Delta>0$ and is very small. $\mathrm{Size}$ of prediction sets corresponding to APS is 3.29. Figure LABEL:fig:mean-size-snaps shows that only a few other labels different from real label have the average scores lower than the quantile of scores. $\mathrm{Size}$ of prediction sets corresponding to SNAPS is 1.29. Overall, for basic non-conformity scores that match this distribution of our assumptions, SNAPS can achieve superior performance based on this scores. The results of CiteSeer and Amazon Computers datasets are available in Appendix C.
Table 2: Ablation study in terms of $\mathrm{Size}$. Overall, three parts of our method are critical and removing any of them results a general decrease in performance.
@@
Ablation study. To understand the effects of three part of our method, i.e., original scores (Orig.), neighborhood scores (Neigh.) and feature similarity node scores (Feat.), we conduct a thorough ablation experiments using GCN at $\alpha=0.05$. In Table 2, SNAPS has the best performance on most of datasets when all three parts are included. Moreover, for remaining dataset on which SNAPS exhibits comparable performance, all those better cases contains the Feat. part. Overall, each part plays a critical role in CP for GNNs and removing any will in general decrease performance.
Parameter analysis. We conduct additional experiments to analyse the robustness of SNAPS. We choose GCN as the GNNs model and APS as the basic non-conformity score function.
Figure LABEL:fig:param-k-size and Figure LABEL:fig:param-k-sh show that as $k$ increases, both $\mathrm{Size}$ and $\mathrm{SH}$ tend to stabilize, because more nodes with the same label are selected to enhance the ego node. On the other hand, we find that when $k$ is extremely large, it appears that nodes with the same label cannot be selected with high accuracy only by feature similarity. Thus, when $k$ is extremely large, performance will decline slightly. Figure LABEL:fig:param-ab-size and Figure LABEL:fig:param-ab-sh show that as the values of parameter $\lambda$ and $\mu$ change, the most areas in the heatmaps of $\mathrm{Size}$ and $\mathrm{SH}$ display similar colors. Overall, SNAPS is robust to the parameter $k$ and is not sensitive to parameter $\lambda$ and $\mu$.
@@
where $s(\boldsymbol{x},y)$ is the score of standard APS, $\mathcal{N}_{\boldsymbol{x}}$ is the $k$ nearest neighbors based on image features in the calibration set and $\eta$ is a corrected weight. We conduct experiments on ImageNet, whose test dataset is equally divided into the calibration set and the test set. For SNAPS, we set $k=5$ and $\eta=0.5$. We report the results of $\mathrm{Coverage}$, $\mathrm{Size}$ and size-stratified coverage violation (SSCV) (Angelopoulos et al., 2021). The details of experiments and SSCV are provided in Appendix D.
@@ 5 Related Work
Uncertainty Quantification for GNNs. Many uncertainty quantification (UQ) methods have been proposed to quantify the model uncertainty for classification tasks in machine learning (Gal and Ghahramani, 2016; Guo et al., 2017b; Zhang et al., 2020; Gupta et al., 2021). Recently, several calibration methods for GNNs have been developed, such as CaGCN (Wang et al., 2021), GATS (Hsu et al., 2022) and SimCalib (Tang et al., 2024). However, these UQ methods lack statistically rigorous and empirically valid coverage guarantee (Huang et al., 2023b). In contrast, SNAPS provides valid coverage guarantee both theoretically and empirically.
Conformal Prediction for GNNs. Although several CP methods for GNNs have been studied, the use of CP in graph-structured data is still largely underexplored. ICP (Wijegunawardana et al., 2020) is the first to apply CP framework on graphs, designs a margin conformity score for labels of nodes without considering the relation between nodes. NAPS (Clarkson, 2023) use the non-exchangeable technique from (Barber et al., 2023) for inductive node classification, not applicable for the transductive setting, while we focus on the transductive setting where exchangeability property holds. Our method is essentially an enhanced version of the DAPS (Zargarbashi et al., 2023) method, which proposes a diffusion-based method that incorporates neighborhood information by leveraging the network homophily. Similar to DAPS, CF-GNN (Huang et al., 2023b) introduces a topology-aware output correction model, akin to GCN, which employs a conformal-aware inefficiency loss to refine predictions and improve efficiency of post-hoc CP. Other recent efforts in conformal prediction for graphs include (Lunde, 2023; Marandon, 2023; Zargarbashi and Bojchevski, 2023; Sanchez-Martin et al., 2024) which focus on distinct problem settings. In this work, SNAPS takes into account both network topology and feature similarity. This method can be applied not only to graph-structure data, but also to other types of data, such as image data.
@@ 6 Conclusion
Our work focuses on node classification using transductive learning. However, in real-world scenarios, many classification tasks require inductive learning. In the future, we aim to apply our method to the inductive setting. Additionally, the method we use to select nodes with the same as the ego node is both computationally inefficient and lacking accuracy. Future work will explore more efficient and accurate methods for node selection. The datasets that we study are not general enough, and we would like to validate our method on more datasets.
@@ A.1 Proof of Proposition 1
Proof. [Zargarbashi et al., 2023] have proved that $(1-\lambda-\mu)\boldsymbol{S}+\mu\boldsymbol{\hat{A}}\boldsymbol{S}$ is exchangeable for $v_{i}\in(\mathcal{V}_{\text{calib}}\cup\mathcal{V}_{\text{test}})$. So we only need to prove that $\boldsymbol{\hat{A}}_{s}\boldsymbol{S}$ is also exchangeable. $\boldsymbol{\hat{A}}_{s}$ is obtained by calculating the feature similarity between two nodes from a global perspective. Before obtaining this matrix, we can not distinguish between labeled and unlabeled nodes, so we just build a new graph structure using node features without considering the order of nodes. So when aggregating non-conformity, we do not break the permutation equivariant. Therefore, $\boldsymbol{\hat{S}}=(1-\lambda-\mu)\boldsymbol{S}+\lambda\boldsymbol{\hat{A}}_{s}\boldsymbol{S}+\mu\boldsymbol{\hat{A}}\boldsymbol{S}$ is a special case of a message passing GNNs layer. It follows that $\boldsymbol{\hat{S}}$ is invariant to permutations of the order of the calibration and testing nodes on the graph. Through the proof above, we can conclude that $\boldsymbol{\hat{S}}$ is also exchangeable for $v_{i}\in(\mathcal{V}_{\text{calib}}\cup\mathcal{V}_{\text{test}})$.
@@ A.2 Proof of Proposition 2
For the sake of description, we denote "$1-\alpha$ quantile of basic non-conformity scores in the calibrated set" as "the quantile score". We first introduce some background before the proof.
Let $\boldsymbol{S}$ and $\boldsymbol{\hat{S}}$ denote APS and SNAPS non-conformity scores, respectively. For node $v$ whose label is $k$, $\boldsymbol{\hat{S}}_{v}$ can be actually be expressed as
@@
In order to prove Proposition 2, we only need to prove the following: 1) SNAPS is efficient for the score corresponding to the ground-truth label $k$ of node $v$, i.e., $\boldsymbol{\hat{S}}_{vk}\leq\boldsymbol{S}_{vk}$ or $\boldsymbol{\hat{S}}_{vk}\leq\eta$. 2) SNAPS is efficient for the score corresponding to the other label $i$ of node $v$, i.e., $\boldsymbol{\hat{S}}_{vi}\geq\boldsymbol{S}_{vi}$ or $\boldsymbol{\hat{S}}_{vi}\geq\eta$. The key idea behind this is as follows. We try to ensure that scores corresponding to the ground-truth label are below the quantile score or decrease compared to the before and scores corresponding to the other label are above the the quantile score or increase compared to the before.
Proof. Here’s our proof.
@@
Thus, $\boldsymbol{\hat{S}}_{vk}<\eta$. This means that for original scores less than the quantile score, they still less than the quantile score after aggregation.
@@
00 0 $<\boldsymbol{S}_{vi}-E_{k}[\boldsymbol{S}_{ui}]$
$\leq\boldsymbol{S}_{vi}-(\eta-\Delta).$
Thus, $\boldsymbol{S}_{vi}-\eta>-\Delta$. Then we have
@@
$=(1-\lambda)(\boldsymbol{S}_{vi}-\eta)+\lambda(E_{k}[\boldsymbol{S}_{ui}]-\eta)$
$>(1-\lambda)(-\Delta)+\lambda(-\Delta)$
$=-\Delta.$
Thus, $\boldsymbol{\hat{S}}_{vi}>\eta-\Delta$. Since $\hat{\eta}<\eta$ and $\Delta$ is very small, the probability of $\boldsymbol{\hat{S}}_{vi}<\hat{\eta}$ is very low. This means even though in this case we may have corrupted some scores corresponding to the other labels, the probability of $\boldsymbol{\hat{S}}_{vi}>\hat{\eta}$ is very high.
@@
$\mathbb{E}[|\mathcal{C}(\boldsymbol{x})|]\leq\mathbb{E}[|\tilde{\mathcal{C}}(\boldsymbol{x})|]$.

### v2

@@ 4.2 Experimental results
SNAPS generates smaller prediction sets and achieves a higher singleton hit ratio. Table 1 shows that $\mathrm{Coverage}$ of all conformal prediction methods is close to the desired coverage $1-\alpha$. At a significance level $\alpha=0.05$, $\mathrm{Size}$ and $\mathrm{SH}$ exhibit superior performance. For example, when evaluated on Products, SNAPS reduces $\mathrm{Size}$ from 14.92 of APS to 7.68. Overall, the experiments show that SNAPS has the desired coverage rate and gets smaller $\mathrm{Size}$ and higher $\mathrm{SH}$ than APS, RAPS, and DAPS. Detailed results for other basic models and SNAPS based on RAPS are available in Appendix D.
SNAPS generates smaller average prediction sets for each label. We conduct additional experiments to analyze the average performance of APS and SNAPS on nodes belonging to the same label at a significance level $\alpha=0.05$. Figure LABEL:fig:mean-size-aps shows that the distribution of the average non-conformity scores for nodes belonging to the same label aligns with the assumptions made in Proposition 2, i.e., $E_{k}[\boldsymbol{S}_{uk}]<\eta$ and $E_{k}[\boldsymbol{S}_{ui}]-\eta\geq-\Delta$, where $\Delta=\eta-(1-\epsilon_{ki})E_{k}[\pi(\boldsymbol{x}_{u})_{max}]-E_{k}[\xi\cdot\pi(\boldsymbol{x}_{u})_{i}]$. If $\Delta>0$, then it is very small. $\mathrm{Size}$ of prediction sets corresponding to APS is 3.29. Figure LABEL:fig:mean-size-snaps shows that only a few other labels different from real labels have average scores lower than the quantile of scores. $\mathrm{Size}$ of prediction sets corresponding to SNAPS is 1.29. Overall, for basic non-conformity scores that match this distribution of our assumptions, SNAPS can achieve superior performance based on these scores. The results of CiteSeer and Amazon Computers datasets are available in Appendix D.
Table 2: Ablation study in terms of $\mathrm{Size}$. Overall, three parts of our method are critical, and removing any of them results in a general decrease in performance.
@@
Ablation study. To understand the effects of three parts of our method, i.e., original scores (Orig.), neighborhood scores (Neigh.), and feature similarity node scores (Feat.), we conduct a thorough ablation experiment using GCN at $\alpha=0.05$. In Table 2, SNAPS performs best on most datasets when all three parts are included. Moreover, for the remaining dataset on which SNAPS exhibits comparable performance, all those better cases contain the Feat. part. Overall, each part plays a critical role in CP for GNNs, and removing any will in general decrease performance.
Table 3: Results of $\mathrm{Coverage}$, $\mathrm{Size}$ and $\mathrm{SH}$ on different datasets. For SNAPS we use the APS score as the basic score and set $\lambda=\mu=1/3$. We report the average calculated from 10 GCN runs with each run of 100 conformal splits at a significance level $\alpha=0.05$. Bold numbers indicate optimal performance.
[…]
Datasets APS RAPS DAPS SNAPS APS RAPS DAPS SNAPS APS RAPS DAPS SNAPS
CoraML 0.950 0.958 0.957 0.951 2.50 2.62 2.32 1.74 43.09 27.34 44.52 54.11
PubMed 0.950 0.968 0.967 0.950 1.82 2.10 2.09 1.61 33.39 14.66 23.27 44.11
CiteSeer 0.951 0.950 0.952 0.950 2.41 2.69 2.16 1.90 48.53 35.37 55.40 58.22
CS 0.950 0.953 0.954 0.950 2.04 1.31 1.33 1.13 64.32 66.91 74.91 85.21
Physics 0.951 0.962 0.962 0.950 1.39 1.44 1.28 1.07 72.44 62.22 77.65 88.58
Computers 0.950 0.950 0.951 0.950 3.01 3.04 2.30 2.01 29.21 9.87 42.19 45.98
Photo 0.949 0.950 0.950 0.950 1.90 1.81 1.56 1.30 54.86 47.27 67.57 79.50
Parameter analysis. We conduct additional experiments to analyze the robustness of SNAPS. We choose GCN as the GNNs model and APS as the basic non-conformity score function.
Figure LABEL:fig:param-k-size and Figure LABEL:fig:param-k-sh demonstrate that the performance of SNAPS significantly improves as $k$ gradually increases from 00. This improvement occurs because the increasing nodes with the same label are selected to enhance the ego node. Subsequently, as $k$ continues to increase, the performance of SNAPS tends to stabilize. On the other hand, we find that when $k$ is extremely large, it appears that nodes with the same label cannot be selected with high accuracy only by feature similarity. Thus, when $k$ is extremely large, performance will decline slightly. Figure LABEL:fig:param-ab-size and Figure LABEL:fig:param-ab-sh show that as the values of parameter $\lambda$ and $\mu$ change, the most areas in the heatmaps of $\mathrm{Size}$ and $\mathrm{SH}$ display similar colors. Overall, SNAPS is robust to the parameter $k$ and is not sensitive to parameters $\lambda$ and $\mu$. To further explore the sensitivity of $\lambda$ and $\mu$ to the performance of SNAPS, we set $\lambda=\mu=1/3$, which indicating that three components of SNAPS are equally weighted. The experimental results in Table 3 demonstrate that SNAPS performs well with these default hyperparameters on most datasets.
@@
where $s(\boldsymbol{x},y)$ is the score of standard APS, $\mathcal{N}_{\boldsymbol{x}}$ is the $k$ nearest neighbors based on image features in the calibration set and $\eta$ is a corrected weight. We conduct experiments on ImageNet, whose test dataset is equally divided into the calibration set and the test set. For SNAPS, we set $k=5$ and $\eta=0.5$. We report the results of $\mathrm{Coverage}$, $\mathrm{Size}$ and size-stratified coverage violation (SSCV) (Angelopoulos et al., 2021). The details of experiments and SSCV are provided in Appendix E.
@@ 5 Related Work
Uncertainty Quantification for GNNs. Many uncertainty quantification (UQ) methods have been proposed to quantify the model uncertainty for classification tasks in machine learning (Gal and Ghahramani, 2016; Guo et al., 2017; Zhang et al., 2020; Gupta et al., 2021). Recently, several calibration methods for GNNs have been developed, such as CaGCN (Wang et al., 2021), GATS (Hsu et al., 2022) and SimCalib (Tang et al., 2024). However, these UQ methods lack statistically rigorous and empirically valid coverage guarantee (Huang et al., 2023b). In contrast, SNAPS provides valid coverage guarantees both theoretically and empirically.
Conformal Prediction for GNNs. Many conformal prediction (CP) methods have been developed to provide valid uncertainty estimates for model predictions in machine learning classification tasks (Romano et al., 2020; Angelopoulos et al., 2021; Liu et al., 2024; Wei and Huang, 2024). Although several CP methods for GNNs have been studied, the use of CP in graph-structured data is still largely underexplored. ICP (Wijegunawardana et al., 2020) is the first to apply CP framework on graphs, designs a margin conformity score for labels of nodes without considering the relation between nodes. NAPS (Clarkson, 2023) use the non-exchangeable technique from (Barber et al., 2023) for inductive node classification, not applicable for the transductive setting, while we focus on the transductive setting where exchangeability property holds. Our method is essentially an enhanced version of the DAPS (Zargarbashi et al., 2023) method, which proposes a diffusion-based method that incorporates neighborhood information by leveraging the network homophily. Similar to DAPS, CF-GNN (Huang et al., 2023b) introduces a topology-aware output correction model, akin to GCN, which employs a conformal-aware inefficiency loss to refine predictions and improve the efficiency of post-hoc CP. Other recent efforts in CP for graphs include (Lunde, 2023; Marandon, 2023; Zargarbashi and Bojchevski, 2023; Sanchez-Martin et al., 2024) which focus on distinct problem settings. In this work, SNAPS takes into account both network topology and feature similarity. This method can be applied not only to graph-structured data but also to other types of data, such as image data.
@@ 6 Conclusion
Our work focuses on node classification using transductive learning. However, in real-world scenarios, many classification tasks require inductive learning. In the future, we aim to apply our method to the inductive setting. Additionally, the method we use to select nodes with the same as the ego node is both computationally inefficient and lacking accuracy. Future work will explore more efficient and accurate methods for node selection. Moreover, while our focus is primarily on datasets with high homophily, many heterophilous networks are prevalent in practice. Consequently, further investigation is essential to enhance the adaptability of SNAPS to these networks.
Acknowledgments
This paper is supported by the National Natural Science Foundation of China (Grant No. 62192783, 62376117), the National Social Science Fund of China (Grant No. 23BJL035), the Science and Technology Major Project of Nanjing (comprehensive category) (Grant No. 202309007), and the Collaborative Innovation Center of Novel Software Technology and Industrialization at Nanjing University.
@@ A.1 Proof of Proposition 1
Proof. [Zargarbashi et al., 2023] have proved that $(1-\lambda-\mu)\boldsymbol{S}+\mu\boldsymbol{\hat{A}}\boldsymbol{S}$ is exchangeable for $v_{i}\in(\mathcal{V}_{\text{calib}}\cup\mathcal{V}_{\text{test}})$. So we only need to prove that $\boldsymbol{\hat{A}}_{s}\boldsymbol{S}$ is also exchangeable for $v_{i}\in(\mathcal{V}_{\text{calib}}\cup\mathcal{V}_{\text{test}})$. $\boldsymbol{\hat{A}}_{s}$ is obtained by calculating the feature similarity between two nodes from a global perspective. Before obtaining this matrix, we can not distinguish between labeled and unlabeled nodes, so we just build a new graph structure using node features without considering the order of nodes. So when aggregating non-conformity, we do not break the permutation equivariant. Therefore, $\boldsymbol{\hat{S}}=(1-\lambda-\mu)\boldsymbol{S}+\lambda\boldsymbol{\hat{A}}_{s}\boldsymbol{S}+\mu\boldsymbol{\hat{A}}\boldsymbol{S}$ is a special case of a message passing GNNs layer. It follows that $\boldsymbol{\hat{S}}$ is invariant to permutations of the order of the calibration and testing nodes on the graph. Through the proof above, we can conclude that $\boldsymbol{\hat{S}}$ is also exchangeable for $v_{i}\in(\mathcal{V}_{\text{calib}}\cup\mathcal{V}_{\text{test}})$.
@@ A.2 Proof of Proposition 2
Lemma 1
As stated in Proposition 2, we have
$E_{k}[\boldsymbol{S}_{ui}]\geq(1-\epsilon_{ki})E_{k}[\pi(\boldsymbol{x}_{u})_{max}]+E_{k}[\xi\cdot\pi(\boldsymbol{x}_{u})_{i}],$
where $E_{k}[\pi(\boldsymbol{x}_{u})_{max}]$ denotes the maximum predicted probability of nodes whose ground-truth labels are $k$, $\epsilon_{ki}$ reflects the model’s error in misclassifying the ground-truth label $k$ as label $i$ and $\xi\in[0,1]$ is a uniformly distributed random variable.
Proof of Lemma 1.
Here, we use APS non-conformity scores as the basic non-conformity scores. Then we have,
$\boldsymbol{S}_{ui}=\sum_{j=1}^{|\mathcal{Y}|}\pi(\boldsymbol{x}_{u})_{j}\mathbb{I}[\pi(\boldsymbol{x}_{u})_{j}>\pi(\boldsymbol{x}_{u})_{i}]+\xi\cdot\pi(\boldsymbol{x}_{u})_{i}.$
Suppose $T$ is the number of nodes whose ground-truth label is label $k$. Below we discuss two cases of $\pi(\boldsymbol{x}_{u})_{i}$:
Case a. If $\pi(\boldsymbol{x}_{u})_{i}$ is the largest predicted probability for node $u$, then $E_{k}[\boldsymbol{S}_{ui}]=E_{k}[\xi\cdot\pi(\boldsymbol{x}_{u})_{i}]=E_{k}[\pi(\boldsymbol{x}_{u})_{max}]+E_{k}[\xi\cdot\pi(\boldsymbol{x}_{u})_{i}]-E_{k}[\pi(\boldsymbol{x}_{u})_{max}]$. Suppose the number of nodes satisfying this case is $A$.
Case b. Otherwise, $E_{k}[\boldsymbol{S}_{ui}]\geq E_{k}[\pi(\boldsymbol{x}_{u})_{max}]+E_{k}[\xi\cdot\pi(\boldsymbol{x}_{u})_{i}]$. Suppose the number of nodes satisfying this case is $B$, where $A+B=T$.
Therefore, summing up $E_{k}[\boldsymbol{S}_{ui}]$ for both cases, we have
$A\cdot E_{k}[\boldsymbol{S}_{ui}]+B\cdot E_{k}[\boldsymbol{S}_{ui}]\geq(A+B)\cdot(E_{k}[\pi(\boldsymbol{x}_{u})_{max}]+E_{k}[\xi\cdot\pi(\boldsymbol{x}_{u})_{i}])-A\cdot E_{k}[\pi(\boldsymbol{x}_{u})_{max}].$
This simplifies to: $E_{k}[\boldsymbol{S}_{ui}]\geq E_{k}[\pi(\boldsymbol{x}_{u})_{max}]+E_{k}[\xi\cdot\pi(\boldsymbol{x}_{u})_{i}]-\frac{A}{T}\cdot E_{k}[\pi(\boldsymbol{x}_{u})_{max}]$.
Let $\epsilon_{ki}=\frac{A}{T}$, which reflects the model’s error in misclassifying the ground-truth label $k$ as label $i$. Therefore, we conclude that: $E_{k}[\boldsymbol{S}_{ui}]\geq(1-\epsilon_{ki})E_{k}[\pi(\boldsymbol{x}_{u})_{max}]+E_{k}[\xi\cdot\pi(\boldsymbol{x}_{u})_{i}]$.
Proof of Proposition 2.
For the sake of description, we denote "$1-\alpha$ quantile of basic non-conformity scores in the calibrated set" as "the quantile score". Let $\boldsymbol{S}$ and $\boldsymbol{\hat{S}}$ denote APS and SNAPS non-conformity scores, respectively. For node $v$ whose label is $k$, $\boldsymbol{\hat{S}}_{v}$ can be be expressed as
@@
In order to prove Proposition 2, we only need to prove the following: 1) SNAPS is efficient for the score corresponding to the ground-truth label $k$ of node $v$, i.e., $\boldsymbol{\hat{S}}_{vk}\leq\boldsymbol{S}_{vk}$ or $\boldsymbol{\hat{S}}_{vk}\leq\eta$. 2) SNAPS is efficient for the score corresponding to the other label $i$ of node $v$, i.e., $\boldsymbol{\hat{S}}_{vi}\geq\boldsymbol{S}_{vi}$ or $\boldsymbol{\hat{S}}_{vi}\geq\eta$. The key idea behind this is as follows. We try to ensure that scores corresponding to the ground-truth label are below the quantile score or decrease compared to the before and scores corresponding to the other label are above the quantile score or increase compared to the before.
@@
Thus, $\boldsymbol{\hat{S}}_{vk}<\eta$. This means that for original scores less than the quantile score, they are still less than the quantile score after aggregation.
@@
$>E_{k}[\boldsymbol{S}_{ui}]-\eta$
$\geq(1-\epsilon_{ki})E_{k}[\pi(\boldsymbol{x}_{u})_{max}]+E_{k}[\xi\cdot\pi(\boldsymbol{x}_{u})_{i}]-\eta.$
Let $\Delta=\eta-(1-\epsilon_{ki})E_{k}[\pi(\boldsymbol{x}_{u})_{max}]-E_{k}[\xi\cdot\pi(\boldsymbol{x}_{u})_{i}]$. If $\Delta\leq 0$, then $\boldsymbol{\hat{S}}_{vi}-\eta>0$. Otherwise, since $\epsilon_{ki}$ reflects the model’s error in misclassifying the ground-truth label $k$ as label $i$, and $\hat{\eta}<\eta$, $\Delta$ is very small, which implies that the probability of $\boldsymbol{\hat{S}}_{vi}<\hat{\eta}$ is very low. Therefore, even though some scores corresponding to other false labels may be corrupted, the probability of $\boldsymbol{\hat{S}}_{vi}>\hat{\eta}$ is very high.
@@
$\mathbb{E}[|\tilde{\mathcal{C}}(\boldsymbol{x})|]\leq\mathbb{E}[|\mathcal{C}(\boldsymbol{x})|]$.